When you are young, the minimum distance your eyes can focus is 20 cm. As you get older, your eyes start to focus objects only at 60 cm. What is the focal length of the lens you need to use to be able to see the objects at 20 cm? You can assume the distance between the eye lens and the retina to be 2 cm.
Answer #1

Using lens formula,

1/f = 1/v - 1/u

1/f = 1/-60 - 1/-20

1/f = - 1/60 + 1/20

Focal length, f = 30 cm

Comment in case any doubt please rate my answer...